OpenStudy (fran345):

Given the functions, f(x) = 3x^2 - 7 and g(x) = x^3 + 1, evaluate f(g(2)).

OpenStudy (jackthegreatest):

first evaluate g(x) = 2^3 + 1 = 9

OpenStudy (jackthegreatest):

f(x) = 3 x 9^2 -7 = 3 x 81 - 7 = 236

OpenStudy (jackthegreatest):

@fran345

OpenStudy (fran345):

thanks